could someone please help me with these fraction.These ones are either multiply or divide:
1) 1 1/2 x 5 2/3
2) 1 5/6 divided by 2
3) 2 1/2 divided by 3 1/3
Thank you for your help in advance I have a quiz coming up and I'm lost on this

Answer by Alan3354(21583)

1) 1 1/2 x 5 2/3
= (3/2)*(17/3) Multiply NUMs and DENs
=(3*17)/(2*3) Cancel the 3's
=17/2, or 8.5 if you prefer
----------------------------
2) 1 5/6 divided by 2
= (11/6)/2
= 11/12
------------
3) 2 1/2 divided by 3 1/3
=(5/2)/(10/3)
Invert the DEN and multiply
=(5/2)*(3/10)
= 15/20
= 3/4